OpenStudy (anonymous):

1 - cos h lim h ->0 --------- = 0 how ? h

OpenStudy (anonymous):

is there anyone to help me.. :) :)

OpenStudy (aravindg):

0

OpenStudy (lgbasallote):

do you know how to use L'Hospital's Rule? And this is a Math question NOT physics

OpenStudy (aravindg):

1-cosx is 2sin^2x/2

OpenStudy (experimentx):

\( 1-\cos x = 1 - \cos^2(x/2) + \sin^2(x/2) = 2 \sin^ (x/2)\) <--- change cosine in to half angle formula

OpenStudy (anonymous):

@AravindG g i know it is 0 but how :) @lgbasallote no i really don't know ..

OpenStudy (lgbasallote):

wow so many possible answers o.O all i know is l'hospital lol

OpenStudy (anonymous):

@experimentX i got u till now further plz :) @lgbasallote it's alright :)

OpenStudy (experimentx):

sinx/ x = 1 <--- make it to this form. since there is square in the top x/2 * sin^2x/2/(x/2)^2 <----now you can solve.

OpenStudy (anonymous):

from where "x/2 * sin^2x/2/(x/2)^2" came ?

OpenStudy (anonymous):

@experimentX from where " x/2 * sin^2x/2/(x/2)^2" came ?

OpenStudy (experimentx):

try balancing things yourself.

OpenStudy (anonymous):

@experimentX i got i it thanks alot :) @AravindG thanks u too :)

OpenStudy (aravindg):

welcomess